There is this homework problem I have for Pre-Algebra. It is on scale factors. This is the problem:
What scale factor was used to build a 55 ft wide billboard from a 25 in. wide model?
I tried to write a proportion (model to actual), and I ended up getting 5 in. to 11 ft. I remember my teacher saying that when doing a scale factor, and it is inches to feet, then it has to be a certain number of inches to 1 ft. He said the exception is that if the scale factor ends up like 1/6 in. to 1 ft because there is no 1/6 inch mark on a ruler. Now, I am confused as to how this problem should be done, and I would be very grateful if someone were to clearly explain how to do this problem.
By the way, I am sorry if the category for this question is incorrect.

Answer by stanbon(75887)

What scale factor was used to build a 55 ft wide billboard from a 25 in. wide model?
-------
Ans: (55*12)in/25in = (11/5)12 = 132/5 = 26 2/5 = 26.4
======
The billboard is 26 2/5 or 132/5 times the model
----
Scale factor: 26 2/5 or 26.4 to 1
===========================
Cheers,
Stan H.